OpenStudy (anonymous):

Which of the following is not a subset of {1, 2, 3}? Ø {0} {1, 2, 3}

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

nothing i need help

jimthompson5910 (jim_thompson5910):

is the number 0 in the list of 1, 2, 3 ?

OpenStudy (anonymous):

\[\left\{ 0 \right\} is \not a \subset of \left\{ 1,2,3 \right\}\] \[subsets of \left\{ 1,2,3 \right\} are\] \[\emptyset,\left\{ 1 \right\},\left\{ 2 \right\},\left\{ 3 \right\}\] \[\left\{ 1,2 \right\},\left\{ 2,3 \right\},\left\{ 1,3 \right\},\left\{ 1,2,3 \right\}\]